OpenStudy (anonymous):

Find the equation of the tangent line to the curve y= (x^2-3)^8 at the point (2,1) using derivatives..

zepdrix (zepdrix):

Buhaha c: Look at that cute lil dawg! c:

zepdrix (zepdrix):

We're looking to form a line which we can either write in slope-intercept form \(\large y=mx+b\) or in point-slope form \(\large y-y_o=m(x-x_o)\). It might make more sense to write it in point-slope form, since we've been given a set of points that we can call \(\large (x_o,y_o)=(2,1)\) Our slope \(\large m\) is given by \(\large m=f'(2)\) So we want to first find the derivative of our function \(\large f\), then plug 2 into it! That will be our \(\large m\).

zepdrix (zepdrix):

Do you need any help finding the derivative of \(\large f(x)\)?

OpenStudy (anonymous):

Ohh, makes so much sense now. I know how to derive it, just not the tangent line and how to find that part, that's all. Thanks!

zepdrix (zepdrix):

Oh ok cool c: